I really need help on this problem please. Given f(x)=x^4 and g(x)=2+3x, find (f*g)(x).

Answer by Theo(13342)

f(x) = x^4
g(x) = 2 + 3x.

find (f*g)(x)

(f*g)(x) = f(x) * g(x) = x^4 * (2 + 3x) = 2x^4 + 3x^5